Murray Motor Company wants you to calculate its cost of common stock. During the next 12 months, the company expects to pay dividends (D1) of $3.70 per share, and the current price of its common stock is $76 per share. The expected growth rate is 7 percent. a. Compute the cost of retained earnings (Ke). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) b. If a $7 flotation cost is involved, compute the cost of new common stock (Kn). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
Answer #1

a.Cost of retained earnings=(D1/Current price)+Growth rate

=(3.7/76)+0.07

=11.87%(Approx).

b.Cost of new common stock=(D1/Current price-Flotation cost)+Growth rate

=(3.7/76-7)+0.07

=(3.7/69)+0.07

=12.36%(Approx).